# Ocmm2)/655361)cm Carbon-Efficient 3D DNN Acceleration: Optimizing Performance and Sustainability

### v1

@@ III-C Die Area Estimation for Embodied Carbon Modeling
The logic die area, on the other hand, is determined primarily by the number of Processing Elements (PEs) in the design. Each PE contains a local buffer and a Multiply-Accumulate (MAC) unit, both of which contribute significantly to the total silicon footprint. The local buffer, implemented as a register file, follows a similar area estimation approach to that of the global SRAM, relying on existing area scaling models. However, the MAC unit is the most area-intensive component of the PE and thus has a significant impact on the overall logic die area.
To further break down the MAC unit’s area composition, we analyze its arithmetic circuits under the widely used bfloat16 representation, which has been extensively adopted in modern DNN accelerators, such as Google’s Tensor Processing Units (TPUs) [26]. Each MAC unit consists of several key arithmetic blocks, including a 7-bit multiplier for mantissa multiplication, two 8-bit adders for exponent addition, and a 24-bit integer adder for accumulating the mantissa results. Among these components, multipliers are by far the most silicon-intensive, dominating the area requirements of the MAC unit. Given their substantial contribution to the overall die size, reducing the area of multipliers directly impacts the total embodied carbon footprint of the accelerator. This observation motivates our focus on employing approximate multipliers, as their adoption can lead to significant reductions in silicon area, thereby lowering the accelerator’s carbon footprint while maintaining computational efficiency.
@@ III-D Approximate multipliers in 3D MAC units
In our design, we replace conventional exact multipliers with area-efficient approximate multipliers from the EvoApprox library [27], which provides a wide range of arithmetic circuits with different trade-offs between computational accuracy and silicon area. Specifically, we target the 7-bit multipliers responsible for mantissa multiplication in the bfloat16 MAC unit. By employing approximate versions of these multipliers, we significantly reduce the logic complexity, leading to lower transistor count and smaller die area. Given that the embodied carbon footprint of a die is proportional to its silicon area, this reduction translates directly into lower emissions during fabrication. To ensure that the use of approximate multipliers does not significantly lower model accuracy, we utilize ApproxTrain [23], a framework designed to systematically simulate the propagation of approximation errors in DNN architectures. ApproxTrain evaluates the impact of replacing exact arithmetic operations with approximate versions, allowing us to quantify the trade-offs between accuracy and area savings.

### v2

@@ III-C Die Area Estimation for Embodied Carbon Modeling
The logic die area, on the other hand, is determined primarily by the number of PEs in the design. Each PE contains a local buffer and a Multiply-Accumulate (MAC) unit, both of which contribute significantly to the total silicon footprint. The local buffer, implemented as a register file, follows a similar area estimation approach to that of the global SRAM, relying on existing area scaling models. However, the MAC unit is the most area-intensive component of the PE and thus has a significant impact on the overall logic die area.
To further break down the MAC unit’s area composition, we analyze its arithmetic circuits under the widely used bfloat16 representation, which has been extensively adopted in modern DNN accelerators, such as Google’s TPUs [26]. Each MAC unit consists of several key arithmetic blocks, including a 7-bit multiplier for mantissa multiplication, two 8-bit adders for exponent addition, and a 24-bit integer adder (23 bits mantissa plus the hidden bit) for accumulating the mantissa results. Among these components, multipliers are by far the most silicon-intensive, dominating the area requirements of the MAC unit. Given their substantial contribution to the overall die size, reducing the area of multipliers directly impacts the total embodied carbon footprint of the accelerator. This observation motivates our focus on employing approximate multipliers, as their adoption can lead to significant reductions in silicon area, thereby lowering the accelerator’s carbon footprint while maintaining computational efficiency.
@@ III-D Approximate multipliers in 3D MAC units
In our design, we replace conventional exact multipliers with area-efficient approximate multipliers from the EvoApprox library [27], which provides a wide range of arithmetic circuits with different trade-offs between computational accuracy and area. Specifically, we target the 7-bit multipliers responsible for mantissa multiplication in the bfloat16 MAC unit. By employing approximate versions of these multipliers, we significantly reduce the logic complexity, leading to lower transistor count and smaller die area. Given that the embodied carbon footprint of a die is proportional to its silicon area, this reduction translates directly into lower emissions. To ensure that the use of approximate multipliers does not significantly lower model accuracy, we utilize ApproxTrain [23], a framework designed to systematically simulate the propagation of approximation errors in DNN architectures. ApproxTrain evaluates the impact of replacing exact arithmetic operations with approximate versions, allowing us to quantify the trade-offs between accuracy and area savings.